Miguel is purchasing pizzas and subs for a large corporate event. Pizzas are $18 each and subs are $11 each. He is buying a total of 20 food items and is spending $304. How many pizzas and how many subs does he need to buy?
Answer #1

Pizza: 18$/piece

Sub: 11$/piece

Let x be the number of Pizzas

Let y be the number of Subs

x+y = 20

Total spend = 304$

18*x + 11*y = 304 -------------------(1)

Multiply by 11 on both sides

11*x + 11*y = 20*11

Subtract the above from 1

7*x = 304-220 =84

x = 12

y = 8

No. of Pizzas is 12 and No. of Subs is 8

Let me know if you need anything else, if not please don't forget to like the answer :)